Find the discriminant. -4n^2+9n-13=-7
Found 2 solutions by fractalier, harpazo:
Answer by fractalier(6550)

Rewriting
-4n^2 + 9n - 13 = -7
we get
-4n^2 + 9n - 6 = 0
The discriminant is b^2 - 4ac. Here that is
9^2 - 4(-4)(-6) = 81 - 96 =
-15

Answer by harpazo(655)

Let d = discriminant
d = b^2 - 4ac
Let a = -4
Let b = 9
Let c = -6
d = (9)^2 -4(-4)(-6)
d = 81 + 16(-6)
d = 81 - 96
d = -15